OpenStudy (anonymous):

Area: A sequence of smaller squares is formed by connecting the midpoints of the sides of a larger square: (a) If the area of the largest square is 1 square unit, determine the first five terms of a sequence that describes the area of each successive square. (b) Use a formula to sum the areas of the first 10 squares.

OpenStudy (anonymous):

I came up with this.... side of square 1 = 1 area = 1 side of square 2 = √2/2 area = 1/2 side of square 3 = 1/2 area of square 3 = 1/4 side of square 4 = √2/4 area of square 4 = 2/16 = 1/8 areas are: 1 , 1/2, 1/4, 1/8 & 1/16 looks like a GS , where a=1, r = 1/2 sum(10) = a(1 - r^10) / (1-r) = 1(1 - (1/2)^(10 )/(1/2) = 2(1 - 1/1024) = 2(1023/1024) = 1023/512